OpenStudy (anonymous):

The volume of a fish tank is 30 cubic feet. If the density is 0.3fish over feet cubed, how many fish are in the tank?

OpenStudy (anonymous):

100 10 9 3

OpenStudy (anonymous):

@paki @Orion1213

OpenStudy (anonymous):

This is a conversion problem based on the ratio. \[\frac{ 0.3 fish }{ ft^3 }\] you are given the density as\[30ft^3\] You can eliminate the variable cubic feet and solve for fish

OpenStudy (anonymous):

Example: \[\frac{ fish }{ ft^3 } \times \frac{ ft^3 }{ 1} = fish\] As you can see the variable cubic feet cancels out if you multiply them with each other

OpenStudy (anonymous):

Leaving you with only fish, the answer you are looking for, so perform the operation and remember that you can also check your work by seeing the resulting units.

OpenStudy (anonymous):

i got the answer 3.3 ...so does that mean the answer would be 3..? @Johnbc

OpenStudy (anonymous):

I think i solved it wrong.. ugh :(

OpenStudy (anonymous):

No check your multiplication.

OpenStudy (anonymous):

Or show how you arrived at your answer.

OpenStudy (anonymous):

i did 0.3/30 * 30/1 I then cross multiplied and got 0.3 over 900 and divided and got 3.3... @Johnbc

OpenStudy (anonymous):

would the answer be 9... And yes I guessed on this answer i am confused.

OpenStudy (anonymous):

Let us try that to see how our units would be.